A company announced a​ "1000 Chips​ Trial," claiming that every​ 18-ounce bag of its cookies contained at least 1000 chocolate chips. Students purchased random bags of cookies from different stores and counted the number of chips in each bag. Some of the data is shown below. Complete parts​ (a) through​ (c) below.

1068

1117

1169

1208

1253

1225

1234

1341

1305

1462

Create a​ 95% confidence interval for the average number of chips.

​(___​,___​) chips
Answer #1

0 solution Given 1068, 1117, 1169, 1808, 1253, 1725, 1234, 1341, 1305 1462 ne 10 T= 1068+1117 +1169 +1208+1253+1225+1234+1341D tamidt = +0.057279 = 226215 C:I = R + talidt * $ = 1238.24 2.26215* 112.9826 10 = 1238.27 80-82262725 (1157-377373, 1319.02